OpenStudy (anonymous):

Prove the Rodrigues formula following the steps below: Let v = x^(n)exp(−x). First show that xdv/dx = (n − x)v. Then derive that xd^(n+2)v/dx^(n+2)+ (1 + x)d^(n+1)v/dx(n+1)+ (n + 1)d^(n)v/dx^n= 0 . (4) Next show that y(x) =exp(x)/n! d^(n)v/dx^(n) is a solution of Laguerre’s differential equation xy"(x)+(1-x)y'(x)+ny=0 using (4)

OpenStudy (amistre64):

Let v = x^(n)exp(−x). First show that x dv/dx = (n − x)v \[v = x^n~e^{−x}\] \[\frac{dv}{dx} = -x^{n}~e^{−x}\] \[x\frac{dv}{dx} = -x^{n+1}~e^{−x}\] \[x\frac{dv}{dx} = (-x)v\] maybe i aint reading this right

OpenStudy (amistre64):

forgot to drop the n

OpenStudy (anonymous):

i have managed the first part i just not sure about the show y(x) is a solution using (4)

OpenStudy (amistre64):

well, when i catch up to you, lets see what i can make of it :) \[v = x^n~e^{−x}\] \[v' = nx^{n-1}~e^{−x}-x^n~e^{−x}\] \[v' = (nx^{n-1}-x^n)~e^{−x}\] \[v' = (nx^{-1}-1)~x^ne^{−x}\] \[xv' = (n-x)v\]

OpenStudy (anonymous):

ok thanks for the help :)... its bugging me aha

OpenStudy (amistre64):

Then derive that \[x \frac{d^{n+2}v}{dx^{n+2}}+ (1 + x)\frac{d^{n+1}v}{dx^{n+1}}+ (n + 1)\frac{d^{n}v}{dx^n} = 0\] \[x v^{(n+2)}+ (1 + x)v^{(n+1)} + (n + 1)v^{(n)} = 0\] ................................. \[xv^{(1)} = (n-x)v^{(0)}\] \[xv^{(n+1)} = (n-x)v^{(n)}\] \[v^{(n+1)}+xv^{(n+2)} = -v^{(n)}+(n-x)v^{(n+1)}\] \[v^{(n+1)}+xv^{(n+2)}+v^{(n)}-(n-x)v^{(n+1)} = 0\] \[xv^{(n+2)}+(x-n+1)v^{(n+1)}+v^{(n)} = 0\] murmur ....

OpenStudy (amistre64):

\[y =\frac{e^x}{n!} v^{(n)}\] \[y^{(1)} =\frac{e^x}{n!} (v^{(n)}+v^{(n+1)})\] \[y^{(2)} =\frac{e^x}{n!} (v^{(n)}+2v^{(n+1)}+v^{(n+2)})\] replace the ys and thier derivatives with these and see if it goes to zero

OpenStudy (amistre64):

assuming of course that i derived those correctly ;)

OpenStudy (anonymous):

ok so the question is essentially saying plug the given y in to laguerres DE?

OpenStudy (amistre64):

yes, and you should be able to do some simplifications with the v parts from the first portion of it

OpenStudy (anonymous):

ok let me try this

OpenStudy (amistre64):

\[xy"+(1-x)y'+ny=0\] \[x(\frac{e^x}{n!} (v^{(n)}+2v^{(n+1)}+v^{(n+2)}))\]\[~~+(1-x)(\frac{e^x}{n!} (v^{(n)}+v^{(n+1)}))\]\[~~~~+n(\frac{e^x}{n!} v^{(n)})=0\] \[\frac{e^x}{n!}\]that factors out and never equals zero so its redundant, leaving us with \[x((v^{(n)}+2v^{(n+1)}+v^{(n+2)}))\]\[~~+(1-x)( (v^{(n)}+v^{(n+1)}))\]\[~~~~+n( v^{(n)})=0\] \[xv^{(n)}+2xv^{(n+1)}+xv^{(n+2)}\]\[~~+(1-x)v^{(n)}+(1-x)v^{(n+1)}\]\[~~~~+n v^{(n)}=0\]

OpenStudy (amistre64):

\[xv^{(n+2)}+(2x+1-x)v^{(n+1)}+(x+1-x+n)v^{(n)}=0\] \[xv^{(n+2)}+(x+1)v^{(n+1)}+(1+n)v^{(n)}=0\]

OpenStudy (anonymous):

Which equals (4) :) you are a lifesaver amistre64. God it was such a basic thing haha. Thanks for the help though much appreciated!!